find tan15 degree using sub multiple formula.solve if u can
 

Grade:10

1 Answers

Avinash
5 years ago
We know that cos∆=1-tan^∆/2//1+tan^∆/2So let ∆=30° and tan^∆/2=XTherefore Cos30°=1-x/1+x√3/2=1-x/1+xCross multiply2-2x=√3+√3x2-√3=x(√3+2);X=2-√3/√3+2Rationalize the denominatorThen we will getX=7-4√3X=4+3-4√3X=[2-√3]^Therefore √X=2-√3Therefore tan∆/2=2-√3Hence proved..